SECOND ORDER REPRESENTATIONS OF THE LEAST ABSOLUTE DEVIATION REGRESSION ESTIMATOR

MIGUEL A. ARCONES

Department of Mathematics, University of Texas, Austin, TX 78712-1082, U.S.A.

(Received April 8, 1996; revised April 28, 1997)

Abstract. We consider exact weak and strong Bahadur-Kiefer representations of the least absolute deviation estimator for the linear regression model. The precise behavior of these representations is obtained under minimal conditions.

Key words and phrases: Least absolute deviation, regression, empirical processes.